# Six loop critical exponent analysis for Lee-Yang and percolation theory

### v1

@@
In recent years there has been a significant advance in the renormalization of core quantum field theories to high loop order due to the development of new techniques such as the Laporta algorithm, [1], the Forcer package, [2, 3], high numerical precision methods based on difference equations to determine massive tadpole graphs, [1], and graphical functions [4, 5]. These tools have opened the door to $\beta$-functions to five loops and beyond [6, 7, 8, 9, 10, 11, 12, 13, 14]. While [6, 7, 8] have extended the renormalization of Quantum Chromodynamics (QCD) to five loops and that of Gross-Neveu-Yukawa theory to the same order, [9], renormalization at orders beyond that have also been achieved with the six loop $\beta$-function of scalar $\phi^{3}$ theory available now in [14] and the seven loop $\phi^{4}$ theory $\beta$-function provided in [11]. Such precision for gauge theories, for instance, has refined uncertainties on observables for particle physics experiments as well as in applications to condensed matter physics based on scalar and Gross-Neveu-Yukawa theories. More specifically for the latter critical exponent estimates governing the properties of phase transitions have been refined.
@@
In previous perturbative analyses of critical exponents in $\phi^{3}$ theory one method of extracting estimates was to use Padé approximants and we follow this approach here. While the Padé method uses rational polynomials to approximate a function in a parameter which is regarded as small there is no guarantee that an approximant will capture the salient features of the exact function. For instance if the function is continuous an approximant may contain discontinuities or if the function is monotonic the Padé may have stationary points. In the present situation the aim is to create approximants that are valid in 22 $\leq$ $d$ $\leq$ 66 and then read off estimates for three, four and five dimensions. The obvious concern is that while $\epsilon$ is regarded as small in relation to the establishment of the Wilson-Fisher fixed point perturbatively it would be unlikely to be applicable when $\epsilon$ == 33. However in previous lower loop analyses the construction of Padé approximants for Lee-Yang and percolation theory exponents benefitted greatly from knowledge of the exact exponents in two dimensions in both instances. These have been established from connecting the critical theory in $d$ $\neq$ 22 to a known conformal field theory in two dimensions which were all classified in [25]. In the Lee-Yang case exact exponents are also available in one dimension. Therefore we will use such available information to construct what is termed two-sided or constrained Padé approximants for each of the above exponents as well as those for percolation.
@@ 22 $\frac{4}{5}$ $\frac{1}{6}$ $\frac{5}{12}$ $\frac{5}{6}$ 22 $\frac{2}{5}$
We have applied this process to all the exponents in (2.6) by first constructing the Padé approximants analytically. Using the numerical representation to select the relevant approximants we arrive at the results of Table 2. For each exponent estimates are recorded in three different dimensions. There are several columns of data. Those headed with $\mu_{L}$ for $L$ == 44, 55 and $6$ are the averages of the selected approximants at $L$ loops only. These are provided as a guide or an indication of the progression of the inclusion of higher order contributions. They are not to be regarded as bona fide estimates. Returning to an earlier point that one can never be sure which approximants will be reliable it turned out that only seven passed the sift test for $\hat{\theta}$ which was the lowest number of all five exponents. In fact no four loop $\hat{\theta}$ ones did which is the reason there are no entries for that exponent. The final two columns contain the two means calculated from all Padés that survived the sifting process where the uncertainties are deduced from the respective $\varsigma$ and $\varsigma_{\mbox{\footnotesize{wt}}}$. There are several general observations. First the uncertainties are invariably smaller as $d$ increases which can be seen across all five exponents in $d$ == 55 with $\hat{\theta}$ being an exception. The reason for this is that there were only seven Padés for $\hat{\theta}$ with the absence of four loop estimates affecting the uncertainty. In places the uncertainties from the weighted measures are tighter than those from the usual mean. This is partly because more approximants were available at five and six loops and their effect can be gauged in the trends apparent in $\mu_{L}$.
@@
While the majority of exponent estimates arise from more modern techniques, such as the functional renormalization group and conformal bootstrap which are continuum field theory based, it is worth discussing our estimates in comparison to those. In particular we concentrate on $\hat{\eta}$, $\hat{\sigma}$ and to a lesser extent on $\hat{\omega}$ as these tend to have been directly measured. What is interesting is that the three dimensional values of $\hat{\eta}$, aside from [36], are very much in accord. In four dimensions a similar picture is apparent but perhaps with much less overlap of uncertainties. In five dimensions the deviation of $\hat{\eta}$ of perturbation theory from the two continuum techniques is more distinct. This seems peculiar in that it might have been expected that the five dimensional perturbative results should be more accurate when summing down from the critical dimension. The much earlier estimates of [26, 33] for $\hat{\eta}$ do not lean towards either side due to the large uncertainties. The position with $\hat{\sigma}$ is not dissimilar unsurprisingly since most methods derive an estimate for $\hat{\eta}$, or $\Delta_{\phi}$ in the case of the conformal bootstrap, and apply the scaling relations (2.7). What is reassuring in the comparison of our results with other approaches is the consistency of the two-sided Padé construction indicating that the resummation down from six dimensions can indeed capture low dimension properties. Finally the situation with our $\hat{\omega}$ is not as conclusive merely because there appears to be only one earlier analysis across all three discrete dimensions, [26]. In that instance our three estimates do not overlap with those of [26]. However comparing our estimate for $\hat{\omega}$ with the recent use of the fuzzy sphere method in three dimensions, [30], is in good agreement.
@@
Recalling that to six loops there is a total of $25$ potential approximants, we have constructed the two-sided Padé approximants with one constraint for each exponent. Of the $25$ we have removed those $d$-dimensional expressions which fail the same the sifting test used for Lee-Yang theory. In the case of $\eta$ no approximants survived. For $\alpha$ and $\Omega$ only nine Padés could be used while for $\omega$ it was eight which was the smallest number for this set of exponents. Indeed this aspect of the analysis is reflected in Table 6 which records our results in a parallel way to Table 2. For instance for these three exponents there was no five loop approximant for $\alpha$ and only one for each of $\omega$ and $\Omega$ which could be used for the means of the final two columns of Table 6. For $\alpha$ there was a degree of stability indicated by the individual four and six loop averages and reflected in the uncertainties. For the two correction to scaling exponents $\Omega$ has a similar stability over the three higher loop means. For $\omega$ this is not the case with the individual loop averages increasing at each loop oder. For the remaining exponents between $12$ and $18$ Padé approximants survived the sieving process with $\delta$, $\sigma$ and $\tau$ being at the upper end.
@@
For the former there is a marked difference in the estimates across the three dimensions in that the central value is less reliable as a guide to behaviour as $d$ decreases. This lies with the fact that the difference in the endpoint boundaries at two and six dimensions, respectively $18.5$ and $2$, is by far the largest range of the exponent set as can be seen in Table 5. So it would be difficult to produce the level of uncertainty of the five dimensional estimate in the three dimensions. To try to address this we followed the same strategy used in [13] which was to construct two-sided Padés for $1/\delta$ in $d$-dimensions and deduce estimates for $\delta$ from the set of valid approximants. In this approach the magnitude of $1/\delta$ only varies by around $0.2$ between the boundary dimensions in contrast to $16.5$ for $\delta$. The associated results for this are recorded against the entry denoted as $\delta^{\dagger}$ in Table 6. While the respective central values are similar to the direct approach for $\delta$ the uncertainties are smaller. The situation with $\sigma$ and $\tau$ is cleaner in that with the endpoint differences being two orders of magnitude less than that of $\delta$ the central values are settled judging by the trend of the four, five and six loop means. This results in significantly smaller uncertainties in a dimensional comparison with $\delta$ with possibly a more accurate three dimensional estimate.

### v2

@@
In recent years there has been a significant advance in the renormalization of core quantum field theories to high loop order due to the development of new techniques such as the Laporta algorithm, [1], the Forcer package, [2, 3], high numerical precision methods based on difference equations to determine massive tadpole graphs, [1], and graphical functions [4, 5]. These tools have opened the door to $\beta$-functions to five loops and beyond [6, 7, 8, 9, 10, 11, 12, 13, 14]. While [6, 7, 8] have extended the renormalization of Quantum Chromodynamics (QCD) to five loops and that of Gross-Neveu-Yukawa theory to the same order, [9], renormalization at orders beyond five have also been achieved with the six loop $\beta$-function of scalar $\phi^{3}$ theory available now in [14] and the seven loop $\phi^{4}$ theory $\beta$-function provided in [11]. Such precision for gauge theories, for instance, has refined uncertainties in observables for particle physics experiments as well as in applications to condensed matter physics based on scalar and Gross-Neveu-Yukawa theories. More specifically for the latter critical exponent estimates governing the properties of phase transitions have been refined.
@@
In previous perturbative analyses of critical exponents in $\phi^{3}$ theory one method of extracting estimates was to use Padé approximants and we follow the same approach here. While the Padé method uses rational polynomials to approximate a function in a parameter which is regarded as small there is no guarantee that an approximant will capture the salient features of the exact function. For instance if the function is continuous an approximant may contain discontinuities or if the function is monotonic the Padé may have stationary points. In the present situation the aim is to create approximants that are valid in 22 $\leq$ $d$ $\leq$ 66 and then read off estimates for three, four and five dimensions. The obvious concern is that while $\epsilon$ is regarded as small in relation to the establishment of the Wilson-Fisher fixed point perturbatively it would be unlikely to be applicable when $\epsilon$ == 33. However in previous lower loop analyses the construction of Padé approximants for Lee-Yang and percolation theory exponents benefitted greatly from knowledge of the exact exponents in two dimensions in both instances. These have been established from connecting the critical theory in $d$ $\neq$ 22 to a known conformal field theory in two dimensions which were all classified in [25]. In the Lee-Yang case exact exponents are also available in one dimension. Therefore we will use such available information to construct what is termed two-sided or constrained Padé approximants for each of the above exponents as well as those for percolation.
@@ 22 $\frac{4}{5}$ $\frac{1}{6}$ $\frac{5}{12}$ $\frac{5}{6}$ 22 $\frac{2}{5}$
We have applied this process to all the exponents in (2.6) by first constructing the Padé approximants analytically. Using the numerical representation to select the relevant approximants we arrive at the results of Table 2. For each exponent estimates are recorded in three different dimensions. There are several columns of data. Those headed with $\mu_{L}$ for $L$ == 44, 55 and $6$ are the averages of the selected approximants at $L$ loops only. These are provided as a guide or an indication of the progression of the inclusion of higher order contributions. They are not to be regarded as bona fide estimates. Returning to an earlier point that one can never be sure which approximants will be reliable it turned out that only seven passed the sift test for $\hat{\theta}$ which was the lowest number of all five exponents. In fact no $\hat{\theta}$ four loop ones did which is the reason there are no entries for that exponent. The final two columns contain the two means calculated from all Padés that survived the sifting process where the uncertainties are deduced from the respective $\varsigma$ and $\varsigma_{\mbox{\footnotesize{wt}}}$. There are several general observations. First the uncertainties are invariably smaller as $d$ increases which can be seen across all five exponents in $d$ == 55 with $\hat{\theta}$ being an exception. The reason for this is that there were only seven Padés for $\hat{\theta}$ with the absence of four loop estimates affecting the uncertainty. In places the uncertainties from the weighted measures are tighter than those from the usual mean. This is partly because more approximants were available at five and six loops and their effect can be gauged in the trends apparent in $\mu_{L}$.
@@
While the majority of exponent estimates arise from more modern techniques, such as the functional renormalization group and conformal bootstrap which are continuum field theory based, it is worth discussing our estimates in comparison to those. In particular we concentrate on $\hat{\eta}$, $\hat{\sigma}$ and to a lesser extent on $\hat{\omega}$ as these tend to have been directly measured. What is interesting is that the three dimensional values of $\hat{\eta}$, aside from [36], are very much in accord. In four dimensions a similar picture is apparent but perhaps with much less overlap of uncertainties. In five dimensions the deviation of $\hat{\eta}$ of perturbation theory from the two continuum techniques is more distinct. This seems peculiar in that it might have been expected that the five dimensional perturbative results should be more accurate when summing down from the critical dimension. The much earlier estimates of [26, 33] for $\hat{\eta}$ do not lean towards either side due to the large uncertainties. The position with $\hat{\sigma}$ is not dissimilar unsurprisingly since most methods derive an estimate for $\hat{\eta}$, or $\Delta_{\phi}$ in the case of the conformal bootstrap, and apply the scaling relations (2.7). What is reassuring in the comparison of our results with other approaches is the consistency of the two-sided Padé construction indicating that the resummation down from six dimensions can indeed capture low dimension properties. Finally the situation with our $\hat{\omega}$ values is not as conclusive merely because there appears to be only one earlier analysis across all three discrete dimensions, [26]. In that instance our three estimates do not overlap with those of [26]. However comparing our estimate for $\hat{\omega}$ with the recent use of the fuzzy sphere method in three dimensions, [30], there is good agreement. By way of a final comment on resummation, an alternative approach to potentially improve convergence of the high order $\epsilon$ series is to use the Padé-Borel method. This involves first replacing the unconstrained series itself by a Borel transform. Its integrand is related to the original series but with the coefficients of $\epsilon$ weighted by $1/L!$ where $L$ is the loop order. The resulting series in the Borel variable is then replaced by a Padé approximant before evaluating the transform for the three values of $\epsilon$ used here. However in following this procedure it is not possible to constrain the integrand approximant by the low dimensional exponent values. Hence the bridge to low dimensions afforded by exact exponent values in one and two dimensions, and of great benefit to the direct Padé approach, cannot be accessed in the Padé-Borel case.
@@
Recalling that to six loops there is a total of $25$ potential approximants, we have constructed the two-sided Padé approximants with one constraint for each exponent. Of the $25$ we have removed those $d$-dimensional expressions which fail the same the sifting test used for Lee-Yang theory. In the case of $\eta$ no approximants survived. For $\alpha$ and $\Omega$ only nine Padés could be used while for $\omega$ it was eight which was the smallest number for this set of exponents. Indeed this aspect of the analysis is reflected in Table 6 which records our results in a parallel way to Table 2. For instance for these three exponents there was no five loop approximant for $\alpha$ and only one for each of $\omega$ and $\Omega$ which could be used for the means of the final two columns of Table 6. For $\alpha$ there was a degree of stability indicated by the individual four and six loop averages and reflected in the uncertainties. For the two correction to scaling exponents $\Omega$ has a similar stability over the three higher loop means. For $\omega$ this is not the case with the individual loop averages increasing at each loop order. For the remaining exponents between $12$ and $18$ Padé approximants survived the sieving process with $\delta$, $\sigma$ and $\tau$ being at the upper end.
@@
For the former there is a marked difference in the estimates across the three dimensions in that the central value is less reliable as a guide to behaviour as $d$ decreases. This lies in the fact that the difference in the endpoint boundaries at two and six dimensions, respectively $18.5$ and $2$, is by far the largest range of the exponent set as can be seen in Table 5. So it would be difficult to produce the level of uncertainty of the five dimensional estimate in the three dimensions. To try to address this we followed the same strategy used in [13] which was to construct two-sided Padés for $1/\delta$ in $d$-dimensions and deduce estimates for $\delta$ from the set of valid approximants. In this approach the magnitude of $1/\delta$ only varies by around $0.2$ between the boundary dimensions in contrast to $16.5$ for $\delta$. The associated results for this are recorded against the entry denoted as $\delta^{\dagger}$ in Table 6. While the respective central values are similar to the direct approach for $\delta$ the uncertainties are smaller. The situation with $\sigma$ and $\tau$ is cleaner in that with the endpoint differences being two orders of magnitude less than that of $\delta$ the central values are settled judging by the trend of the four, five and six loop means. This results in significantly smaller uncertainties in a dimensional comparison with $\delta$ with possibly a more accurate three dimensional estimate.